OpenStudy (ria23):

For f(x)=1/x-2 and g(x)=x^2+3 a.) (f o g)(x) b.) (g o f)(8) The work I did for this was wrong, can someone walk me through how to do this right?

OpenStudy (jhannybean):

a) \(f(g(x)) \implies \ f(x^2+3)\) so \[f(g(x))=\frac{1}{(x^2+3)-2}\]

OpenStudy (jhannybean):

b) g(f(8)) --> \(\left(\frac{1}{x-2}\right)^2 +3\)

OpenStudy (jhannybean):

For (a) all you have to do is simplify the denominator, and you'll have your answer :)

OpenStudy (jhannybean):

for (b) you have \(\left(\frac{1}{x-2}\right)^2 \implies \frac{1}{(x-2)^2} \implies \frac{1}{x^2 -4x+4} \)

OpenStudy (ria23):

Thank yhu! c: I can see where I went wrong. Thank yhu so much

OpenStudy (jhannybean):

No problem :)